OpenStudy (anonymous):

The local skating rink pays Mary a fixed rate per pupil plus a base amount to work as a skating instructor. She earns $90 for instructing 15 students on Monday afternoon. Last Friday, she earned $62 for working with 8 students. Lisa is also a skating instructor. She receives half the base amount that Mary does, but she is paid twice as much per student. Who would earn more money instructing a class of 20 students?

OpenStudy (texaschic101):

Mary -- f = fixed rate per person b = base amount b + 15f = 90 --- multiply by -1 b + 8f = 62 ----------- -b - 15f = -90 (result of multiplying by -1) b + 8f = 62 ------------add -7f = - 28 f = -28/-7 f = 4 b + 15f = 90 b + 15(4) = 90 b + 60 = 90 b = 90 - 60 b = 30 Mary earns $30 base pay and charges $4 per person Lisa earns half base pay of Mary.....30/2 = 15 paid twice as much per student ... 4 x 2 = 8 so Lisa earns $15 base pay and charges $8 per person =============================== for 20 students... Mary makes : b + 20h = total earned for 20 students 30 + 20(4) = 30 + 80 = 110 so Mary makes $110 for teaching 20 students Lisa makes : b + 20(8) = total earned for 20 students 15 + 160 = 175 so Lisa makes $175 for teaching 20 students ================================ so who earns more ??